OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5. _x__y=_?

OpenStudy (anonymous):

use y=mx+b plug the points given in to y and x appropriately to find b. as well as slope in for m. then set y=(3/5)=b, and find whatever you got for b into the equation and thats your answer.

OpenStudy (anonymous):

lil confused

OpenStudy (anonymous):

sorry. well i dont want to just give you the answer so i will help you get the answers yourself. first, a general equation of a line is y=mx+b. where m is slope, b is the y intercept, and y and x are self explainitory. the first thing you are going to is plug your two given points, and slope, into the equation for x and y appropriately. so you equation will look like: -4=2(3/5)+b. now can you solve for b?

OpenStudy (anonymous):

okay! this is makin sense lol

OpenStudy (anonymous):

haha ok. so once you solve for b, all you do is plug your slope and b back into the equation. so you final equation will look like: y=(3/5)x+b and just plug whatever you found b to be, into the equation. and it is simple as that.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem!